OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation?

OpenStudy (anonymous):

x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

@ParthKohli please help:)

Parth (parthkohli):

If \(|x| = a \), then \(x = a ~\text{or}~-a\)

OpenStudy (anonymous):

i dont understand @ParthKohli

Parth (parthkohli):

Hmm ... for example, if \(|y| = 4\), then \(y\) is either 4 or -4.

OpenStudy (anonymous):

-x - 2 = 4 i got this @ParthKohli